# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ 6 Post-Training Certification
As mentioned previously, there is an appeal in having ML model architectures that are easily integrable in post-training verification frameworks. We now showcase how to integrate SCNNs in such frameworks by proposing a local-Lipschitz stability certification program to evaluate the stability induced by different SCNN training programs, viz., (WDR1-SCNNT), (WDR2-SCNNT), and (SCNNT). We certify the worst-case output deviation of SCNNs when subject to a small perturbation $\varepsilon\in\mathcal{E}$ in their feature space. This methodology is inspired by Huang et al. (2021) on $\operatorname{ReLU}$ feedforward deep neural networks, but tailored to SCNNs thanks to the decoupled formulation. Recall the definition of the Lipschitz constant for a function $f:\mathcal{X}\to\mathbb{R}$. A function is $K$-Lipschitz if there exists a constant $0<K<+\infty$ such that :
$\frac{|f(\mathbf{x}_{1})-f(\mathbf{x}_{2})|}{|\mathbf{x}_{1}-\mathbf{x}_{2}|}\leq K,$
for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$. In our case, we consider only neighboring points by defining $\mathbf{x}_{2}=\{\mathbf{x}_{1}+\bm{\varepsilon}|\bm{\varepsilon}\in\mathcal{E},\mathbf{x}_{1}\in\mathcal{X},\mathcal{E}\text{ is a small convex set}\}$, thus the term local-Lipschitz. We define the $\varepsilon$-stability of a function $f$ by $L_{\varepsilon}$ such that:
$L_{\varepsilon}=\max_{\mathbf{x}_{1},\bm{\varepsilon}:\mathbf{x}_{1}\in\mathcal{X},\bm{\varepsilon}\in\mathcal{E}}|f(\mathbf{x}_{1})-f(\mathbf{x}_{1}+\bm{\varepsilon})|,$
which is visibly linked to the local-Lipschitz bound.
@@
Let $\bm{\beta}^{\star}$ be the optimal weights obtained by an SCNN training program and a set of sampling vectors $\mathcal{S}$ obtained before training. The worst possible output deviation $0<L_{\varepsilon}<+\infty$ of the SCNN for any perturbation vector $\bm{\varepsilon}$, from a small convex set $\mathcal{E}$, on any input from the input space can be obtained by a mixed-integer convex program (MICP).
Proof: Let $M\in\mathbb{R}_{>0}$ a large strictly positive scalar, $\mathbf{x}\in\mathcal{X}$ any possible input from feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$, $\mathbf{x}_{\varepsilon}\in\mathcal{X}$ a perturbed input close to $\mathbf{x}$. The post-training verification problem can be formulated as:
@@
$\mathbf{x}$ $\in{\Xi},$ (31)
$\mathbf{x},{\mathbf{x}}_{\varepsilon}$ $\in\mathcal{X}$ (32)
where $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{P})$ and $\mathbf{B}=(\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{P})$ are intermediate variables. $\mathcal{E},\Xi\subset\mathbb{R}^{d}$ are convex sets that ensure that $\bm{\varepsilon}$ is finite and small and that $\mathbf{x}$ is finite, respectively, e.g., box constraints on each element or bounded Euclidian norms. Constraints (10)--(18) and (20)--(28) are used to represent the $\operatorname{ReLU}$-like activation patterns of the SCNN for the unperturbed and perturbed output, respectively. Specifically, matrices $\mathbf{A}$ and $\mathbf{B}$ model $\mathbf{xd}^{\top}$ and $\mathbf{x_{\varepsilon}d_{\varepsilon}}^{\top}$ with constraints (14)--(17) and (24)--(27), respectively. Constraints (12), (13), (22), and (23) embed the following definitions: $\mathbf{d}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$ and $\mathbf{d}_{\varepsilon}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$. Finally, the non-concave objective function can be linearized using another binary variable:
$\max_{\hat{y},\hat{y}_{\varepsilon},\mathbf{x},\bm{\varepsilon},\mathbf{A},\mathbf{B},\bm{\gamma},\bm{\eta}}\alpha$ (33)
@@
$\square$
Besides pure certification, this formulation is interesting because it could be used in a pessimistic bilevel program, see Dempe (2002), that integrates SCNN training and certification into the same problem. We leave this opening for future work. We remark that $M$ can be tuned separately for each constraint to be as tight as possible and that recent works have greatly accelerated similar procedures for deep $\operatorname{ReLU}$ neural networks, e.g., see Zhou et al. (2024).
@@ 7 Outlier generation on benchmark functions
In this paper, to generate empirical distributions with complex patterns, we propose to use the mathematical expressions of traditional non-convex benchmark functions as a starting point. These functions are normally used to evaluate optimization solvers because of their interesting patterns and their generalizability to $n$ dimensions.
@@
We use a methodology inspired by the work of Pallage et al. (2024) to generate outliers. They propose a method based on the sliced-Wasserstein (SW) distance to remove out-of-distribution outlier points from training sets. In this work, we do the opposite and use a SW-based method to ensure that introduced outliers are out-of-distribution with respect to the original data distribution.
@@
Computing the Wasserstein distance is generally of high computational complexity, but the one-dimensional Wasserstein distance is an exception as it possesses a closed-form solution. The SW distance exploits this special case by first computing, through linear projections, an infinite amount of one-dimensional representations for each original $n$-dimensional distribution. It then computes the average Wasserstein distance between their one-dimensional projections (Kolouri et al. 2019). In practice, the order-$t$ SW distance’s approximation under a Monte Carlo scheme of $L$ samples is defined as:
@@
$SW_{\|\cdot\|,t}\left(\mathbb{U},\mathbb{V}\right)\approx\left(\frac{1}{L}\sum_{l=1}^{L}W_{\|\cdot\|,t}\left(\mathfrak{R}\mathbb{U}\left(\cdot,\theta_{l}\right),\mathfrak{R}\mathbb{V}\left(\cdot,\theta_{l}\right)\right)^{t}\right)^{1/t},$
@@
where $\mathfrak{R}\mathbb{D}(\cdot,\bm{\theta}_{l})$ is a single projection of the Radon transform of distribution function $\mathbb{D}$ over a fixed sample $\bm{\theta}_{l}\in\mathbb{S}^{d-1}=\{\bm{\theta}\in\mathbb{R}^{d}|\theta_{1}^{2}+\theta_{2}^{2}+\cdots+\theta_{d}^{2}=1\}$ $\forall l\in\llbracket L\rrbracket$. Note that $\mathbb{S}^{d-1}$ is often referred to as the $(d-1)$-dimensional unit sphere as it encloses every $d$-dimensional unit-norm vector in a sphere of radius 1. It can be shown that the WD and SWD are closely related (Bonnotte 2013, Theorem 5.1.5). We refer interested readers to Bonneel et al. (2015) and Kolouri et al. (2019) as they offer deeper mathematical insights.
@@
Our procedure is as follows. Consider an empirical distribution of $N$ samples ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$ obtained by sampling uniformally a non-convex benchmark function denoted by $\mathbb{P}$. We use the notation ^$\hat{\mathbb{P}}_{N+m}$ to denote the empirical distribution with $m$ added samples from $\mathbb{P}$ and ^$\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m}$ to represent the union of the empirical distribution with a set of $m$ outlier samples $\mathcal{O}_{m}$. Let $\varphi_{m}>0$ represent a threshold SW distance for sample $m$. The SW distance deals with distributions of equal weights and must account for present outliers when generating a new one. Assuming we want to generate a total of $\overline{m}\in\mathbb{Z}_{>0}$ out-of-sample outliers, the iterative procedure is based on:
@@
^$\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m}=\left\{\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\}\ \middle|\tilde{\mathbf{z}}\in\mathcal{Z},\left(SW_{\|\cdot\|,t}(\hat{\mathbb{P}}_{N+1}\cup\mathcal{O}_{m-1},\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\})\geq\varphi_{m}\right)\right\}\quad m\in\llbracket\overline{m}\rrbracket,$
@@
where $\mathcal{O}_{0}=\emptyset$ is the empty set. This procedure ensures that each new outlier creates a shift of at least $\varphi$ with respect to the empirical distribution and previous outliers 1, $11,\ 2,\ldots,m-$ 1. We must consider previous outliers to ensure we do not create any new involuntary pattern in the data. Because we do not make a single assumption on $\mathbb{P}$, this methodology is adaptable to generate outliers on any function.
@@ 8.2.1 Setting
In this experiment, we seek to predict the year-round hourly energy consumption in kWh of 16 distinct non-residential buildings located in three adjacent distribution substations in Montréal, Québec, Canada. As we do not have any information on building activity, this task is done using only hourly consumption history and historical meteorological features. We use a 12-week history to train each model to predict the following week. We have approximately 2.5 years of collected data on each building. We proceed with a 12-week sliding window and collect the predictions for each new week. We use the data of the first year and a half for hyperparameter tuning and the data of the last year for performance assessment. The features and label are listed in Table 2.
@@ 1 Hourly energy consumption [kWh] $\mathbb{R}_{+}$
We set cyclical features with a $\cos/\sin$ encoding similarly to (Moon et al. 2018) and introduce a binary flag to indicate weekends and public holidays. The dataset is created from proprietary data obtained through our collaboration with Hydro-Québec.

### v3

@@ 6 Post-Training Certification
As mentioned in Section 1, there is value in having ML model architectures that are easily integrable in post-training verification frameworks. Indeed, PTV enhances trustworthiness in the ML pipeline by providing theoretical certifications for trained models. In this section, we demonstrate how to integrate SCNNs in such frameworks by proposing a local-Lipschitz-like stability certification program to evaluate the stability induced by different SCNN training programs, viz., (WDR1-SCNNT), (WDR2-SCNNT), and (SCNNT). Let $\mathcal{E}\subset\mathbb{R}^{d}$ be a convex bounded set of input perturbations. Our goal is to certify the worst-case output deviation of trained SCNNs when subject to a small bounded perturbation $\varepsilon\in\mathcal{E}$ in their feature space in a way that is agnostic to the training method. This will allow us to compare the stability induced by different regularization paradigms in Section 8.3 without mapping the SCNN weights to their non-convex equivalent. This methodology is inspired by the work of Huang et al. (2021) on $\operatorname{ReLU}$ feedforward deep neural networks, but tailored to SCNNs thanks to the decoupled formulation.
Recall the definition of the Lipschitz constant for a function $f:\mathcal{X}\to\mathbb{R}$. A function is $K$-Lipschitz if there exists a constant $0<K<+\infty$ such that :
${|f(\mathbf{x}_{1})-f(\mathbf{x}_{2})|}\leq K{\|\mathbf{x}_{1}-\mathbf{x}_{2}\|},$
for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$. In our case, we consider only neighbouring points by defining $\mathbf{x}_{2}=\{(\mathbf{x}_{1}+\bm{\varepsilon}){{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ \in\mathcal{X}}}\ |\ \bm{\varepsilon}\in\mathcal{E},\mathbf{x}_{1}\in\mathcal{X}\}$, thus the term local-Lipschitz. We define the $\varepsilon$-stability of a function $f$ by $L_{\varepsilon}$ such that:
$L_{\varepsilon}=\max_{\mathbf{x}_{1},\bm{\varepsilon}:\mathbf{x}_{1}\in\mathcal{X},\bm{\varepsilon}\in\mathcal{E},\mathbf{x}_{1}+\bm{\varepsilon}\in\mathcal{X}}|f(\mathbf{x}_{1})-f(\mathbf{x}_{1}+\bm{\varepsilon})|,$
which is visibly linked to the local-Lipschitz bound as $L_{\varepsilon}$ englobes the product of $K$ and $\|\bm{\varepsilon}\|$, given the definition of $\mathbf{x}_{2}$ in the neighbourhood of $\mathbf{x}_{1}$.
@@
Let $\bm{\beta}^{\star}$ be the optimal weights obtained by an SCNN training program and a set of sampling vectors $\mathcal{S}$ obtained before training. Assuming an SCNN is a finite mapping from a bounded feature space to the label space, the worst possible output deviation $0\leq L_{\varepsilon}<+\infty$ of the SCNN for any perturbation vector $\bm{\varepsilon}$, from a bounded convex set $\mathcal{E}$, on any input from a bounded convex feature space $\mathcal{X}$ can be obtained by a mixed-integer convex program (MICP).
Proof: Recall the formulation of the modified samples from Example 1. Let $M\in\mathbb{R}_{>0}$ a large positive scalar, i.e., a big-$M$ (Williams 1978), $\mathbf{x}\in\mathcal{X}$ any possible input from feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$, $\mathbf{x}_{\varepsilon}\in\mathcal{X}$ a perturbed input in the neighbourhood of $\mathbf{x}$. The post-training verification problem can be formulated as:
@@
$\mathbf{x},{\mathbf{x}}_{\varepsilon}\$$\in\mathcal{X}$ (31)
where $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{P})$ and $\mathbf{B}=(\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{P})$ are intermediate variables used to represent modified samples in their matrix form before vectorization, see Example 1. Here, $\mathcal{E}\subset\mathbb{R}^{d}$ is a bounded convex set that defines the neighbourhood around a given sample $\mathbf{x}$ from the feature space $\mathcal{X}$, e.g., box constraints on each element or bounded Euclidean norm. The magnitude of $\mathcal{E}$ should be chosen with respect to the application.
Constraints (10)-(18) and (20)-(28) are used to represent the $\operatorname{ReLU}$-like activation patterns of the SCNN for the unperturbed and perturbed output, respectively. Specifically, matrices $\mathbf{A}$ and $\mathbf{B}$ model the outer products $\mathbf{xd}^{\top}$ and $\mathbf{x_{\varepsilon}d_{\varepsilon}}^{\top}$ with constraints (14)-(17) and (24)-(27), respectively. These constraints ensure that if the product $\mathbf{x}^{\top}\mathbf{s}_{i}$ is positive, i.e., $\gamma_{i}=1$, then $a_{ki}=x_{k},\ \forall k\in\llbracket d\rrbracket$. Otherwise, if $\gamma_{i}=0$, then we have $a_{ki}=0,\ \forall k\in\llbracket d\rrbracket$. The process is similar for the product $\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{i}$ but with matrix $\mathbf{B}$ and binary variable $\bm{\eta}$, instead. Constraints (12), (13), (22), and (23) embed the following definitions: $\mathbf{d}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$ and $\mathbf{d}_{\varepsilon}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$, given a known set of sampling vectors $\mathcal{S}$. In sum, we use the binary variables $\bm{\gamma}$ and $\bm{\eta}$ to model the result of the indicator function for samples $\mathbf{x}$ and $\mathbf{x}_{\varepsilon}$, respectively. Finally, the non-concave objective function can be linearized using another binary variable, $\zeta$, and disjunctive constraints, yielding:
@@
In (32)-(39), we once again use a binary variable to ensure that if $\hat{y}_{\varepsilon}>\hat{y}$, then $\alpha=\hat{y}_{\varepsilon}-\hat{y}$; if $\hat{y}_{\varepsilon}<\hat{y}$, then $\alpha=\hat{y}-\hat{y}_{\varepsilon}$; and if $\hat{y}_{\varepsilon}=\hat{y}$, then $\alpha=0$. Thus, $\alpha$ is equivalent to the absolute value objective. This concludes the proof. $\square$
Besides pure certification, this formulation is interesting because it could be used in a pessimistic bilevel program, see Dempe (2002), that integrates SCNN training and certification into the same problem. We leave this idea as a direction for future research. We remark that $M$, as in any big-$M$ formulation used to model disjunctive constraints, can be tuned separately for each constraint to be as tight as possible such that each disjunctive constraint holds. Recent works have greatly accelerated similar procedures for deep $\operatorname{ReLU}$ neural networks, e.g., see Zhou et al. (2024).
@@ 7 Outlier generation on benchmark functions
In this section, we outline the methodology used to generate outlier-filled synthetic datasets for the numerical evaluation of WaDiRo-SCNNs in Section 8.1. The proposed methodology is original as it leverages an optimal transport-based criterion for outlier generation and does not make any assumption on the data-generating function, allowing for a wide range of possible synthetic datasets.
@@
First, to generate empirical distributions with complex patterns, we propose to use the mathematical expressions of traditional non-convex benchmark functions as a starting point. These functions are normally used to evaluate optimization solvers because of their interesting patterns and their generalizability to $n$ dimensions.
@@
We use a methodology inspired by the work of Pallage et al. (2024) and Pallage and Lesage-Landry (2025) to generate outliers. They propose a method based on the sliced-Wasserstein (SW) distance to remove out-of-distribution outlier points from training sets. In this work, we do the opposite and use a SW-based method to ensure that introduced outliers are out-of-distribution with respect to the original data distribution.
@@
Computing the Wasserstein distance is generally of high computational complexity, but the one-dimensional Wasserstein distance is an exception as it possesses a closed-form solution. The SW distance (SWD) exploits this special case by first computing, through linear projections, an infinite amount of one-dimensional representations for each original $n$-dimensional distribution. It then computes the average Wasserstein distance between their one-dimensional projections (Kolouri et al. 2019). In practice, the order-$t$ SW distance’s approximation under a Monte Carlo scheme of $L$ samples is defined as:
@@
$SW_{\|\cdot\|,t}\left(\text{U},\text{V}\right)\approx\left(\frac{1}{L}\sum_{l=1}^{L}W_{\|\cdot\|,t}\left(\mathfrak{R}\text{U}\left(\cdot,\theta_{l}\right),\mathfrak{R}\text{V}\left(\cdot,\theta_{l}\right)\right)^{t}\right)^{1/t},$
@@
where $\mathfrak{R}\text{D}(\cdot,\bm{\theta}_{l})$ is a single projection of the Radon transform of distribution function D over a fixed sample $\bm{\theta}_{l}\in\mathbb{S}^{d-1}=\{\bm{\theta}\in\mathbb{R}^{d}|\theta_{1}^{2}+\theta_{2}^{2}+\cdots+\theta_{d}^{2}=1\}$ $\forall l\in\llbracket L\rrbracket$. Note that $\mathbb{S}^{d-1}$ is often referred to as the (d1)(d-1)-dimensional unit sphere as it encloses every $d$-dimensional unit-norm vector in a sphere of radius 1. It can be shown that the WD and SWD are closely related (Bonnotte 2013, Theorem 5.1.5). We refer interested readers to Bonneel et al. (2015) and Kolouri et al. (2019) as they offer deeper mathematical insights.
@@
Our procedure is as follows. Consider an empirical distribution of $N$ samples $\hat{\text{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$ obtained by sampling uniformally a non-convex benchmark function on the domain of interest. This domain is predetermined. Each sample is constructed as a feature and label pair where the features are composed of the uniform samples and the labels are the respective evaluation of the benchmark function on each feature. We denote the data-generating distribution as P. We use the notation $\hat{\text{P}}_{N+m}$ to denote the empirical distribution with $m$ added samples from P and $\hat{\text{P}}_{N}\cup\mathcal{O}_{m}$ to represent the union of the empirical distribution with a set of $m$ outlier samples $\mathcal{O}_{m}$. Let $\varphi_{m}>0$ represent a threshold SW distance for sample $m$. The SW distance deals with distributions of equal weights and must account for present outliers when generating a new one. Assuming we want to generate a total of $\overline{m}\in\mathbb{N}_{>0}$ out-of-sample outliers, the iterative procedure is based on:
@@
$\hat{\text{P}}_{N}\cup\mathcal{O}_{m}=\left\{\hat{\text{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\}\ \middle|\tilde{\mathbf{z}}\in\mathcal{Z},\left(SW_{\|\cdot\|,t}(\hat{\text{P}}_{N+1}\cup\mathcal{O}_{m-1},\hat{\text{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\})\geq\varphi_{m}\right)\right\}\ m\in\llbracket\overline{m}\rrbracket,$
@@
where $\mathcal{O}_{0}=\emptyset$ is the empty set. This procedure ensures that each new outlier creates a shift of at least $\varphi$ with respect to the empirical distribution and previous outliers 1, $1,\ 2,\ldots,m$-1. We must consider previous outliers to avoid creating any new involuntary patterns in the data. Because we do not make a single assumption on P, this methodology is adaptable to generate outliers on any function. This will allow us to test empirically the distributional robustness of different machine learning models against out-of-distribution outliers, with respect to the SW distance, on synthetic datasets with rich patterns.
@@ 8.2.1 Setting
In this experiment, we seek to predict the year-round hourly energy consumption in kWh of 16 distinct non-residential buildings located in three adjacent distribution substations in Montréal, Québec, Canada. These buildings are enrolled in a non-residential DR program. Being able to predict the baseline consumption of these assets under different contexts allows VPPs to better characterize grid needs through time and to retroactively estimate the relief given by each building during DR events. As we do not have any information on building activity, this task is done using only hourly consumption history and historical meteorological features. We use a 12-week history to train each model to predict the following week. We have approximately 2.5 years of collected data on each building. We proceed with a 12-week sliding window and collect the predictions for each new week. We use the data of the first year and a half for hyperparameter tuning and the data of the last year for performance assessment. The features and label are listed in Table 2.
@@ 1 Hourly energy consumption [kWh] $\mathbb{R}_{+}$
We set cyclical features with a $\cos/\sin$ encoding similarly to (Moon et al. 2018) and introduce a binary flag to indicate weekends and public holidays. The cyclical encoding consists of projecting each feature space onto the unit circle to model proximity between lower and higher values of a temporal feature, e.g., 23:00 is close to 1:00, but this is not represented numerically as it is. The dataset is created from proprietary data obtained through our collaboration with Hydro-Québec.